OpenStudy (anonymous):

find x and y so that abcd is a parallelogram AB= 3x+2y BC=3 CD=7 DA= 2x+y

OpenStudy (mathstudent55):

In a parallelogram, opposite sides are congruent. Then AB = CD and BC = DA 3x + 2y = 7 2x + y = 3 Multiply the second equation by -2 and add to equation 1: 3x + 2y = 7 -4x - 2y = - 6 --------------(add) -x = 1 x = -1 Now substitute x = -1 in first original equation: 3x + 2y = 7 3(-1) + 2y = 7 -3 + 2y = 7 2y = 10 y = 5